Use Principle of Mathematical Induction to show that for all n e N, an = 212.521 11 + 32 .221 11 is divisible by 19.
Answer #1

Let P(n) be the given statement that is P(n): an = 2+2.52n + 4 ght? 2n+1 is divisible by 19 when n=r PU: Q1 = 25+3?2? = 103mf2 2.2 52.5.2m +1 3.3992 2² 22m ft = 50.2me? 52m +1 + 12.3m+? 2 2m + 1 mt2 = 12 (2m+? 52m++ 37*222*7) + 38 mai f mf2 amfil